# Unravelling the emergence of quantum jumps in a monitored qubit

### v1

@@ A Postselected Liouvillian with Decoherence
The first two transitions are determined by the no-click dynamics, governed by the postselected Liouvillian $\tilde{\mathcal{L}}_{P}$ of equation (S26). To analyse its eigenvalue structure, we parametrise the density matrix in the $\left|1\right\rangle$$\left|0\right\rangle$ basis as
@@ C The Second Transition: Onset of State Freezing
and the condition $e_{2}=2e_{3}$ yields $\lambda^{\mathrm{c}}_{2}=2/\sqrt{3}\approx 1.155$. With finite decoherence, however, the transition location shifts (Fig. S4). Notably, dephasing lowers the location of the second transition, in contrast to its effect on the first transition. Moreover, false positives do impact the second transition, increasing $\lambda^{\mathrm{c}}_{2}$. However, in our experiment, Schmitt-trigger filtering (Supplementary Information Section 3) strongly suppresses the effective false-positive rate, well below its raw value $\kappa_{\mathrm{FP}}\equiv p_{\mathrm{FP}}/dt\approx 125\,\mathrm{ms}^{-1}$. Therefore, we set $\kappa_{\mathrm{FP}}=0$ in the realistic model. For our experimental parameters, numerical analysis yields $\lambda^{\mathrm{c}}_{2}\approx 1.01$, confirming that dephasing can invert the ordering of the first two transitions.
@@ D The Third Transition: Onset of Quantum Zeno Dynamics
The ensemble dynamics including the detector’s excited state are given by equation (S27). Extending the parametrisation (S28) to include the detector population, we define $\bm{v}=[p_{B},p_{S},x,z]^{T}$. The evolution becomes $d\bm{v}/dt=\mathcal{L}_{\mathrm{B}}\bm{v}$, with
@@
The third transition occurs when the characteristic polynomial $\mathcal{C}_{B}(\Lambda)=\det(\mathcal{L}_{B}-\Lambda\mathbb{I})$ has a repeated root. The waiting time $\tau_{\mathrm{B}}$ interpolates between two limits: for $\tau_{\mathrm{B}}\to 0$, the system returns instantly from $\left|\mathrm{B}\right\rangle$ to $\left|0\right\rangle$, recovering the standard ensemble dynamics with $\lambda^{\mathrm{c}}_{3}=2$ in the ideal case. For $\tau_{\mathrm{B}}\to\infty$, the system never returns, and the ensemble dynamics reduce to the postselected dynamics, giving $\lambda^{\mathrm{c}}_{3}\to\lambda^{\mathrm{c}}_{1}$. As shown in Fig. S4(e), $\lambda^{\mathrm{c}}_{3}$ drops significantly even for modest $\tau_{\mathrm{B}}$: for $\kappa\equiv\Omega_{\mathrm{S}}\tau_{\mathrm{B}}=2.5$ (our experimental value), numerical diagonalisation gives $\lambda^{\mathrm{c}}_{3}\approx 1.25$.
SI 7 Extraction of Observed Transition Values
@@ A First transition — $\lambda^{\mathrm{obs}}_{1}$
Figure S5: Extracting $\lambda^{\mathrm{obs}}_{1}$ from the pole spectrum of the no-click probability $P^{(0)}(t)$. a, Real parts of the three poles governing $P^{(0)}(t)$; b, c, imaginary and real components of the pole splitting $\Delta e_{12}=(e_{1}-e_{2})/2$, all obtained from the matrix pencil method. Points show the bootstrap median and error bars indicate a $1\sigma$-equivalent uncertainty. The transition is characterised by the vanishing of the imaginary component and the emergence of a real component. The dashed black line shows a joint fit of the real and imaginary parts to equation (3) near the critical point. The vertical red line marks the location of the first transition, measured to be $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$.
@@
We extract the three poles of $P^{(0)}(t)$ using the matrix pencil method [35], a high-resolution technique for estimating the poles of complex exponentials, setting the model order to three. Confidence intervals are estimated with a residual bootstrap procedure separately for each $\lambda$. First, a fitted curve is constructed from the poles and amplitudes of the complex exponentials obtained with the matrix pencil method. Then, residuals are calculated by subtracting the fit from the experimental $P^{(0)}(t)$ data. In each of the $2000$ bootstrap iterations, synthetic data are generated by sampling these residuals with replacement and adding them to the fitted curve. The matrix pencil method is then reapplied to the synthetic data. Consistent tracking of the pole labelling across different bootstrap iterations is enforced by a fixed sorting convention. Finally, we record the median and $1\sigma$ confidence interval for both the real and imaginary components of each pole, see Fig. S5.
@@
At the critical measurement strength $\lambda^{\mathrm{obs}}_{1}$, two eigenvalues coalesce, and the Liouvillian in equation (S30) acquires a Jordan block structure. Expanding around this transition, the eigenvalue splitting depends on the perturbation $\delta=\lambda-\lambda^{\mathrm{obs}}_{1}$. While the individual eigenvalues also experience common-mode shifts, these cancel out when taking the difference between them. Consequently, the splitting itself is governed by the expansion of the discriminant of the characteristic polynomial, $\sqrt{\alpha\delta+\beta\delta^{2}+\dots}$. Taylor expanding this term yields an expansion of half-integer powers: $\delta^{1/2},\delta^{3/2},\dots$ [42], justifying the form of equation (3), which captures the critical point and the behaviour up to second order.
@@
By treating the two components $\mathrm{Re}(\Delta e_{12})$ and $\mathrm{Im}(\Delta e_{12})$ as a combined data set, we fit them simultaneously to equation (3) with a single shared parameter set. Since the expansion holds locally, the fit only uses a narrow window around the initial guess for $\lambda^{\mathrm{obs}}_{1}$. Varying the size of this window over a reasonable range leaves $\lambda^{\mathrm{obs}}_{1}$ unchanged within uncertainties, yielding $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$.
@@
Benchmarking this extraction procedure on simulations of $P^{(0)}(t)$ including shot noise indicates a systematic underestimate of the extracted $\lambda^{\mathrm{obs}}_{1}$ by $\approx 0.1$. This bias may partially account for the discrepancy between the realistic model (Supplementary Information Section 5) and the observed value.
@@ B Second transition — $\lambda^{\mathrm{obs}}_{2}$
where we treat $(A,\theta_{+},\xi)$ as free parameters. We fit the experimental dwell-time data for each $\lambda$ around an initially guessed $\theta_{+}$ to the bin-averaged model of equation S34:
@@
We then fit the extracted $\xi(\lambda)$ to equation 5, allowing a horizontal shift $\delta\lambda$:
@@
Figure S6: Extraction of $\lambda^{\mathrm{obs}}_{2}$ from the fitted critical exponent $\xi$. This figure complements Fig. 4 by showing the fitted values of $\xi$ obtained from the angular dwell-time data $\tau_{\theta}(\theta)$. Each $\tau_{\theta}(\theta)$ trace is fitted near its first non-vanishing bin ($\approx\theta_{+}$) to the analytical model described in Methods, yielding the parameters $\xi$ (top), $\theta_{+}$ (bottom), and a scaling factor AA (not shown). The dashed black line shows the fit of $\xi(\lambda)$ to the analytical model. The second transition is identified from this fit at $\lambda^{\mathrm{obs}}_{2}=0.92\pm 0.01$ (dashed red vertical line), corresponding to the zero crossing $\xi(\lambda^{\mathrm{obs}}_{2})=0$ (horizontal grey dashed line), where the dwell time transitions from finite to divergent.
@@ C Third transition — $\lambda^{\mathrm{obs}}_{3}$
The third transition is extracted from the spectral analysis of the ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$, modelled by the Liouvillian in equation (S33). Although the full spectrum contains four eigenvalues, for our experimental parameters one real eigenvalue is large and negative. The corresponding eigenmode decays too rapidly to be resolved, given the experimental time resolution $T_{\text{int}}$, rendering it effectively unobservable. Furthermore, one eigenvalue is strictly zero, representing a steady-state solution. The remaining two eigenvalues form a complex-conjugate pair. This effective three-pole structure is confirmed by the singular value decomposition (SVD) step of the matrix pencil method [35], which reveals a rank three model — a zero pole and a complex-conjugate pair. We extract all three poles and track the two oscillatory poles, $e_{1}$ and $e_{2}$. Analogous to the analysis of $\lambda^{\mathrm{obs}}_{1}$, we identify the transition at the coalescence point of this pair. We jointly fit both components of their difference $\mathrm{Re}(\Delta e_{12})$ and $\mathrm{Im}(\Delta e_{12})$ to the near-transition form of equation (3) with a shared fitting parameter set, yielding $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$ (Fig. S7).
Figure S7: Extracting $\lambda^{\mathrm{obs}}_{3}$ from the pole spectrum of the ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$. a, Real parts of the three poles governing the dynamics of $P_{\left|1\right\rangle}^{\mathrm{ens}}$;b, c, imaginary and real components of the pole splitting $\Delta e_{12}=(e_{1}-e_{2})/2$. Points show the bootstrap median and error bars indicate a $1\sigma$-equivalent uncertainty. As in the first transition, the critical point is identified by the simultaneous vanishing of the imaginary component and the emergence of a real component. The dashed black line shows a joint fit of the real and imaginary parts to equation (3). The vertical red line marks the location of the third transition, measured at $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$.
Similar to the first transition, the extraction procedure may underestimate $\lambda^{\mathrm{obs}}_{3}$, potentially accounting in part for the discrepancy with the prediction from the realistic model (Supplementary Information Section 5).
SI 8 Numerical Simulations and Error Budget
@@
To emulate the experimental detection process, we group the simulation into intervals of duration $T_{\mathrm{int}}$} (corresponding to $\frac{T_{\mathrm{int}}}{dt_{\mathrm{sim}}}$ time steps) and compute the total time $T_{\mathrm{B}}$ spent in $\left|\mathrm{B}\right\rangle$ during each interval. If $T_{\mathrm{B}}\geq T_{\text{int}}/2$, the interval is classified as a click; otherwise, it is classified as no-click. Detection errors are then applied by flipping the click (no-click) outcome with probability $p_{\mathrm{FN}}$ ($p_{\mathrm{FP}}$), as obtained from the HMM analysis described in Section 3. Note that a brief excursion to $\left|\mathrm{B}\right\rangle$ lasting less than $T_{\text{int}}/2$ results in a missed detection even before the detection error is applied. Fig. S8 presents simulation results corresponding to the experimental data in Fig. 3. Panel a shows the simulated qubit $\mathrm{Z}$-polarisation conditioned on no clicks. Panel b displays the no-click probability $P^{(0)}(\lambda,t)$ and panel c summarises the error budget by showing the fraction of trajectories affected by each error type.
Figure S8: Numerical simulation of the no-click evolution. a, Simulated qubit $\mathrm{Z}$-polarisation as a function of the no-click sequence duration and measurement strength $\lambda$. Each pixel corresponds to the average of $3\times 10^{8}$ stochastic trajectories, though the effective number decreases at longer durations where a click typically occurs during a trajectory (see b). The simulated behaviour is in good agreement with the experimental data (cf. Fig. 3a). The horizontal grey lines indicate the $\lambda$-values of the one-dimensional cuts in Fig. 3b — one below the first transition ($\lambda=0.5$) and one above it ($\lambda=2.2$). b, probability of obtaining no clicks up to time $t$, $P^{(0)}(\lambda,t)$. Most trajectories experience a click before reaching the target duration. c, Error budget. Each panel shows the fraction of trajectories that reached the target duration but experienced an event of the type indicated.
Figure S9: Cryogenic microwave wiring diagram. The experimental device is enclosed within OFHC copper and A4K shields for thermal and magnetic isolation. Control signals are synthesised by a Quantum Machines OPX+ system and upconverted using IQ-mixers before entering the cryostat. They pass through a sequence of attenuators, low-pass filters (LPF), and infrared filters to ensure proper thermalisation and noise suppression. The output signal is first amplified by a travelling-wave parametric amplifier (TWPA, Silent Waves Dreadnought) and then routed through a double-junction isolator, enabling high-fidelity single-shot readout.

### v2

@@ A Postselected Liouvillian with Decoherence
The first two transitions are determined by the no-click dynamics, governed by the postselected Liouvillian $\tilde{\mathcal{L}}_{P}$ of equation (S29). To analyze its eigenvalue structure, we parametrize the density matrix in the $\left|1\right\rangle$$\left|0\right\rangle$ basis as
@@ C The Second Transition: Onset of State Freezing
and the condition $e_{2}=2e_{3}$ yields $\lambda^{\mathrm{c}}_{2}=2/\sqrt{3}\approx 1.155$. With finite decoherence, however, the transition location shifts (Fig. S4). Notably, dephasing lowers the location of the second transition, in contrast to its effect on the first transition. Moreover, false positives do impact the second transition, increasing $\lambda^{\mathrm{c}}_{2}$. However, in our experiment, Schmitt-trigger filtering (supplementary text section 3) strongly suppresses the effective false-positive rate, well below its raw value $\kappa_{\mathrm{FP}}\equiv p_{\mathrm{FP}}/dt\approx 125\,\mathrm{ms}^{-1}$. Therefore, we set $\kappa_{\mathrm{FP}}=0$ in the realistic model. For our experimental parameters, numerical analysis yields $\lambda^{\mathrm{c}}_{2}\approx 1.01$, confirming that dephasing can invert the ordering of the first two transitions.
@@ D The Third Transition: Onset of Quantum Zeno Dynamics
The ensemble dynamics including the detector’s excited state are given by equation (S30). Extending the parametrization (S31) to include the detector population, we define $\bm{v}=[p_{B},p_{S},x,z]^{T}$. The evolution becomes $d\bm{v}/dt=\mathcal{L}_{\mathrm{B}}\bm{v}$, with
@@
The third transition occurs when the characteristic polynomial $\mathcal{C}_{B}(\Lambda)=\det(\mathcal{L}_{B}-\Lambda\mathbb{I})$ has a repeated root. The waiting time $\tau_{\mathrm{B}}$ interpolates between two limits: for $\tau_{\mathrm{B}}\to 0$, the system returns instantly from $\left|\mathrm{B}\right\rangle$ to $\left|0\right\rangle$, recovering the standard ensemble dynamics with $\lambda^{\mathrm{c}}_{3}=2$ in the ideal case. For $\tau_{\mathrm{B}}\to\infty$, the system never returns, and the ensemble dynamics reduce to the postselected dynamics, giving $\lambda^{\mathrm{c}}_{3}\to\lambda^{\mathrm{c}}_{1}$. As shown in Fig. S4E, $\lambda^{\mathrm{c}}_{3}$ drops significantly even for modest $\tau_{\mathrm{B}}$: for $\kappa\equiv\Omega_{\mathrm{S}}\tau_{\mathrm{B}}=2.5$ (our experimental value), numerical diagonalization gives $\lambda^{\mathrm{c}}_{3}\approx 1.25$.
SM 7 Extraction of Observed Transition Values
@@ A First transition — $\lambda^{\mathrm{obs}}_{1}$
Figure S5: Extracting $\lambda^{\mathrm{obs}}_{1}$ from the pole spectrum of the no-click probability $P^{(0)}(t)$. (A) Real parts of the three poles governing $P^{(0)}(t)$; (B), (C), imaginary and real components of the pole splitting $\Delta e_{12}=(e_{1}-e_{2})/2$, all obtained from the matrix pencil method. Points show the bootstrap median and error bars indicate a $1\sigma$-equivalent uncertainty. The transition is characterized by the vanishing of the imaginary component and the emergence of a real component. The dashed black line shows a joint fit of the real and imaginary parts to equation (S1) near the critical point. The vertical red line marks the location of the first transition, measured to be $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$.
@@
We extract the three poles of $P^{(0)}(t)$ using the matrix pencil method [37], a high-resolution technique for estimating the poles of complex exponentials, setting the model order to three. Confidence intervals are estimated with a residual bootstrap procedure separately for each $\lambda$. First, a fitted curve is constructed from the poles and amplitudes of the complex exponentials obtained with the matrix pencil method. Then, residuals are calculated by subtracting the fit from the experimental $P^{(0)}(t)$ data. In each of the $2000$ bootstrap iterations, synthetic data are generated by sampling these residuals with replacement and adding them to the fitted curve. The matrix pencil method is then reapplied to the synthetic data. Consistent tracking of the pole labeling across different bootstrap iterations is enforced by a fixed sorting convention. Finally, we record the median and $1\sigma$ confidence interval for both the real and imaginary components of each pole, see Fig. S5.
@@
At the critical measurement strength $\lambda^{\mathrm{obs}}_{1}$, two eigenvalues coalesce, and the Liouvillian in equation (S33) acquires a Jordan block structure. Expanding around this transition, the eigenvalue splitting depends on the perturbation $\delta=\lambda-\lambda^{\mathrm{obs}}_{1}$. While the individual eigenvalues also experience common-mode shifts, these cancel out when taking the difference between them. Consequently, the splitting itself is governed by the expansion of the discriminant of the characteristic polynomial, $\sqrt{\alpha\delta+\beta\delta^{2}+\dots}$. Taylor expanding this term yields an expansion of half-integer powers: $\delta^{1/2},\delta^{3/2},\dots$ [44], justifying the form of equation (S1), which captures the critical point and the behavior up to second order.
@@
By treating the two components $\mathrm{Re}(\Delta e_{12})$ and $\mathrm{Im}(\Delta e_{12})$ as a combined data set, we fit them simultaneously to equation (S1) with a single shared parameter set. Since the expansion holds locally, the fit only uses a narrow window around the initial guess for $\lambda^{\mathrm{obs}}_{1}$. Varying the size of this window over a reasonable range leaves $\lambda^{\mathrm{obs}}_{1}$ unchanged within uncertainties, yielding $\lambda^{\mathrm{obs}}_{1}=0.99\pm 0.01$.
@@
Benchmarking this extraction procedure on simulations of $P^{(0)}(t)$ including shot noise indicates a systematic underestimate of the extracted $\lambda^{\mathrm{obs}}_{1}$ by $\approx 0.1$. This bias may partially account for the discrepancy between the realistic model (supplementary text section 5) and the observed value.
@@ B Second transition — $\lambda^{\mathrm{obs}}_{2}$
where we treat $(A,\theta_{+},\xi)$ as free parameters. We fit the experimental dwell-time data for each $\lambda$ around an initially guessed $\theta_{+}$ to the bin-averaged model of equation (S37):
@@
We then fit the extracted $\xi(\lambda)$ to equation (S3), allowing a horizontal shift $\delta\lambda$:
@@
Figure S6: Extraction of $\lambda^{\mathrm{obs}}_{2}$ from the fitted critical exponent $\xi$. This figure complements Fig. 4 by showing the fitted values of $\xi$ obtained from the angular dwell-time data $\tau_{\theta}(\theta)$. Each $\tau_{\theta}(\theta)$ trace is fitted near its first non-vanishing bin ($\approx\theta_{+}$) to the analytical model described in Methods, yielding the parameters $\xi$ (top), $\theta_{+}$ (bottom), and a scaling factor AA (not shown). The dashed black line shows the fit of $\xi(\lambda)$ to the analytical model. The second transition is identified from this fit at $\lambda^{\mathrm{obs}}_{2}=0.92\pm 0.01$ (dashed red vertical line), corresponding to the zero crossing $\xi(\lambda^{\mathrm{obs}}_{2})=0$ (horizontal gray dashed line), where the dwell time transitions from finite to divergent.
@@ C Third transition — $\lambda^{\mathrm{obs}}_{3}$
The third transition is extracted from the spectral analysis of the ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$, modeled by the Liouvillian in equation (S36). Although the full spectrum contains four eigenvalues, for our experimental parameters one real eigenvalue is large and negative. The corresponding eigenmode decays too rapidly to be resolved, given the experimental time resolution $T_{\text{int}}$, rendering it effectively unobservable. Furthermore, one eigenvalue is strictly zero, representing a steady-state solution. The remaining two eigenvalues form a complex-conjugate pair. This effective three-pole structure is confirmed by the singular value decomposition (SVD) step of the matrix pencil method [37], which reveals a rank three model — a zero pole and a complex-conjugate pair. We extract all three poles and track the two oscillatory poles, $e_{1}$ and $e_{2}$. Analogous to the analysis of $\lambda^{\mathrm{obs}}_{1}$, we identify the transition at the coalescence point of this pair. We jointly fit both components of their difference $\mathrm{Re}(\Delta e_{12})$ and $\mathrm{Im}(\Delta e_{12})$ to the near-transition form of equation (S1) with a shared fitting parameter set, yielding $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$ (Fig. S7).
Figure S7: Extracting $\lambda^{\mathrm{obs}}_{3}$ from the pole spectrum of the ensemble-averaged excited state probability $P_{\left|1\right\rangle}^{\mathrm{ens}}$. (A) Real parts of the three poles governing the dynamics of $P_{\left|1\right\rangle}^{\mathrm{ens}}$; (B), (C) imaginary and real components of the pole splitting $\Delta e_{12}=(e_{1}-e_{2})/2$. Points show the bootstrap median and error bars indicate a $1\sigma$-equivalent uncertainty. As in the first transition, the critical point is identified by the simultaneous vanishing of the imaginary component and the emergence of a real component. The dashed black line shows a joint fit of the real and imaginary parts to equation (S1). The vertical red line marks the location of the third transition, measured at $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$.
Similar to the first transition, the extraction procedure may underestimate $\lambda^{\mathrm{obs}}_{3}$, potentially accounting in part for the discrepancy with the prediction from the realistic model (supplementary text section 5).
SM 8 Numerical Simulations and Error Budget
@@
To emulate the experimental detection process, we group the simulation into intervals of duration $T_{\mathrm{int}}$} (corresponding to $\frac{T_{\mathrm{int}}}{dt_{\mathrm{sim}}}$ time steps) and compute the total time $T_{\mathrm{B}}$ spent in $\left|\mathrm{B}\right\rangle$ during each interval. If $T_{\mathrm{B}}\geq T_{\text{int}}/2$, the interval is classified as a click; otherwise, it is classified as no-click. Detection errors are then applied by flipping the click (no-click) outcome with probability $p_{\mathrm{FN}}$ ($p_{\mathrm{FP}}$), as obtained from the HMM analysis described in Section 3. Note that a brief excursion to $\left|\mathrm{B}\right\rangle$ lasting less than $T_{\text{int}}/2$ results in a missed detection even before the detection error is applied. Fig. S8 presents simulation results corresponding to the experimental data in Fig. 3. Panel A shows the simulated qubit $\mathrm{Z}$-polarization conditioned on no clicks. Panel B displays the no-click probability $P^{(0)}(\lambda,t)$ and panel C summarizes the error budget by showing the fraction of trajectories affected by each error type.
Figure S8: Numerical simulation of the no-click evolution. (A) Simulated excited state population as a function of the no-click sequence duration and measurement strength $\lambda$. Each pixel corresponds to the average of $3\times 10^{8}$ stochastic trajectories, though the effective number decreases at longer durations where a click typically occurs during a trajectory (see panel B). The simulated behavior is in good agreement with the experimental data (cf. Fig. 3A). The horizontal gray lines indicate the $\lambda$-values of the one-dimensional cuts in Fig. 3B — one below the first transition ($\lambda=0.5$) and one above it ($\lambda=2.2$). (B) probability of obtaining no clicks up to time $t$, $P^{(0)}(\lambda,t)$. Most trajectories experience a click before reaching the target duration. (C) Error budget. Each panel shows the fraction of trajectories that reached the target duration but experienced an event of the type indicated.
Figure S9: Cryogenic microwave wiring diagram. The experimental device is enclosed within OFHC copper and A4K shields for thermal and magnetic isolation. Control signals are synthesized by a Quantum Machines OPX+ system and upconverted using IQ-mixers before entering the cryostat. They pass through a sequence of attenuators, low-pass filters (LPF), and infrared filters to ensure proper thermalization and noise suppression. The output signal is first amplified by a traveling-wave parametric amplifier (TWPA, Silent Waves Dreadnought) and then routed through a double-junction isolator, enabling high-fidelity single-shot readout.